# LLM-Text Watermarking based on Lagrange Interpolation

### v2

@@ 7 Experiments
Given a short text, such as a paragraph consisting of 300 tokens, we can embed a 300-bit watermark. In practice, a key challenge lies in selecting the appropriate dimension of the Galois field used for the computation. If we choose $GF(2^{6})$, each point occupies 6 bits. This means the text can accommodate 50 watermark points. According to Table 4, only 6 of these points need to survive for the correct watermark line to be recovered, while the remaining 44 points may be replaced by fake or random ones without affecting recoverability. If we instead use $GF(2^{8})$, the total number of watermark points decreases to 36. In this case, as shown in Table 4, just 4 genuine points are enough to reconstruct the correct line, while the remaining 32 random points form lines with a maximum of only 3 points. With $GF(2^{12})$, the number of embedable points drops further to 25. Still, only 3 genuine watermark points are needed to reconstruct the original line, as 22 random points only form trivial two-point lines. A broader discussion considering different text lengths – such as a paragraph, a page, and a 10-page report – is provided in Table 4. It is evident that for longer documents, using Galois fields of higher dimension is advantageous, as the appearance of a three-point line becomes strong evidence of the watermark’s presence and authenticity.
Document Number of tokens Arithmetics Total number Number of Number of
(bits) GF of points watermark points fake points
$GF(2^{6})$ 50 6 44
Paragraph 300 $GF(2^{8})$ 36 4 32
$GF(2^{12})$ 25 3 22
$GF(2^{6})$ 116 7 109
Page 700 $GF(2^{8})$ 87 5 82
$GF(2^{12})$ 58 4 54
$GF(2^{6})$ 1166 26 1140
Report 7,000 $GF(2^{8})$ 875 10 865
(10 pages) $GF(2^{12})$ 583 5 578
$GF(2^{16})$ 437 3 433

### v3

@@ 7 Experiments
Given a short paragraph consisting of 100 tokens. We can embed a 100-bit watermark. In practice, a key challenge lies in selecting the appropriate dimension of the Galois field used for the computation. If we choose $GF(2^{6})$, each point occupies 6 bits. This means the text can accommodate 16 watermark points. According to Table 4, only 8 of these points need to survive for the correct watermark line to be recovered, while the remaining 75 points may be replaced by fake or random ones without affecting recoverability. If we instead use $GF(2^{8})$, the total number of watermark points decreases to 12. In this case, as shown in Table 4, just 6 genuine points are enough to reconstruct the correct line, while the remaining 77 random points form lines with a maximum of only 5 points. With $GF(2^{12})$, the number of watermark points drops further to 8. Still, only 4 genuine watermark points are needed to reconstruct the original line, as 77 random points with a maximum of only 3 points. A broader discussion considering different text lengths – such as a short paragraph, a paragraph, and a 1-page report – is provided in Table 4. It is evident that for longer documents, using Galois fields of higher dimension is advantageous, as the appearance of a three-point line becomes strong evidence of the watermark’s presence and authenticity.
Document # of Tokens Arithmetics Total # of Points # of Watermark # of Fake # of Points Needed
(bits) GF $n(N-1)+1$ Points $N$ Points to Recover $K$
Short $GF(2^{6})$ 91 16 75 8
Paragraph 100 $GF(2^{8})$ 89 12 77 6
$GF(2^{12})$ 85 8 77 4
$GF(2^{6})$ 295 50 245 12
Paragraph 300 $GF(2^{8})$ 289 37 252 6
$GF(2^{12})$ 289 25 264 4
$GF(2^{6})$ 691 116 575 18
Report 700 $GF(2^{8})$ 689 87 602 10
(1 page) $GF(2^{12})$ 685 58 627 4
$GF(2^{16})$ 673 43 630 3